OpenStudy (anonymous):

NO CLUE HOW TO DO THIS. Given that (-9,3) is a point on the graph of f(x), find the corresponding point of the function f(-3x)

OpenStudy (anonymous):

helppppp

OpenStudy (unklerhaukus):

y = f(x)

OpenStudy (anonymous):

well i am on another problem and the points are 5,-4 and on the graph of f(x) and its saying find the corresponding points for -1/4 f(x). So would i start by putting -4=1/4f(x) ?

OpenStudy (unklerhaukus):

the two points on the line are (-9,3) & (5,-4) is that right?

OpenStudy (anonymous):

no those are two separate problems. I couldn't figure out the other one so i guessed and now am on this one: Given that (5,-4) is a point on the graph of f(x), find the corresponding point of the function -1/4f(x).

OpenStudy (anonymous):

it only gives one point on the line, 5,-4

OpenStudy (unklerhaukus):

hmm, is it -1 / ( 4f(x) ), or -(1/4)f(x).

OpenStudy (anonymous):

-(1/4)f(x)

OpenStudy (unklerhaukus):

ok, so (x,y) is (5,-4) y = f(x) substitute to find the value of -(1/4)f(x)

OpenStudy (anonymous):

-4=f(5) ???

OpenStudy (unklerhaukus):

well, that ^ is not false, but it's not what i mean, what is y equal to at the point?

OpenStudy (anonymous):

-4

OpenStudy (unklerhaukus):

yes so sub this into -(1/4)f(x)

OpenStudy (anonymous):

-(1/4)f(-4)

OpenStudy (unklerhaukus):

like this... -(1/4)[-4] the just multiply them

OpenStudy (anonymous):

1

OpenStudy (unklerhaukus):

Yeah!

OpenStudy (anonymous):

so then what do i do from there one point is 1 how do i find the other just substitute 5 in to f(x)?

OpenStudy (unklerhaukus):

you have the same x=5 -(1/4)f(5) = 1

OpenStudy (unklerhaukus):

the point (x,y) is ... (5,1)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

but how do you get 5(-1/4)=1 ??

OpenStudy (unklerhaukus):